6. Prove the following relation:
Answer #1

We have, H = U + PV

or dH = dU+PdV+VdP = TdS+Vdp

partial derivatives , (with V at constant T )

(Invalid EquationH / Invalid EquationV)T = T (Invalid EquationS/Invalid EquationV) T + V (Invalid EquationP/Invalid EquationV) T

or (Invalid EquationH / Invalid EquationV)T = T (Invalid EquationP/Invalid EquationT) V + V (Invalid EquationP/Invalid EquationV) T   (maxwell relation (Invalid EquationS/Invalid EquationV) T = (Invalid EquationP/Invalid EquationT) V )\

Euler theorem ;  (Invalid EquationP/Invalid EquationT) V (Invalid EquationT/Invalid EquationV) P (Invalid EquationV/Invalid EquationP) T  = -1

Using this,   (Invalid EquationH / Invalid EquationV)T = T (Invalid EquationP/Invalid EquationT) V - V  (Invalid EquationP/Invalid EquationT) P (Invalid EquationT/Invalid EquationV) V

or   (Invalid EquationH / Invalid EquationV)T = (Invalid EquationP/Invalid EquationT) V ( T- V  ( (Invalid EquationT/Invalid EquationV) P ) .............................i

We solve, L.H.S.

on multiplying by, V2/V2

we get    V2 (Invalid EquationP/Invalid EquationT) V ( T/V2 - (1/V)* ( (Invalid EquationT/Invalid EquationV) P )) ..................................ii

we have ,   (Invalid Equation(T/V) /Invalid EquationV) P =  (1/V)* ( (Invalid EquationT/Invalid EquationV) P ) -  T/V2 ............................iii

From i, ii, and iii

we get ,   (Invalid EquationH / Invalid EquationV)T = - V2 (Invalid EquationP/Invalid EquationT) V   (Invalid Equation(T/V) /Invalid EquationV) P